40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 5 09:20:44 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_Chamber_optical_layout.pdf
    Reply  Thu Aug 5 12:39:56 2021, Aidan, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber Cryo_in-air_setup.pdf
       Reply  Thu Aug 5 14:29:42 2021, Stephen, Laser, 2um Photodiodes, Optical design for 2um PD in new cryo chamber IMG_9685.JPGIMG_9687.JPGrouting_markup_of_IMG_9688.pngIMG_9682.JPG
          Reply  Tue Aug 24 15:01:25 2021, Stephen, Laser, 2um Photodiodes, 2um PD in new cryo chamber 
             Reply  Thu Aug 26 14:31:46 2021, Aidan, Stephen, Laser, 2um Photodiodes, Incident power calibration of 2um PD in new cryo chamber IMG_4420.jpgREF_PD_POW_CAL_cryo.pdf
                Reply  Fri Aug 27 17:57:30 2021, Stephen, DailyProgress, 2um Photodiodes, Start of PD pumpdown  
                   Reply  Sat Aug 28 07:01:21 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  20210828_-_IMG_4445.jpg20210828_-_IMG_4443.jpg
                      Reply  Sun Aug 29 11:17:43 2021, Aidan, DailyProgress, 2um Photodiodes, Start of PD pumpdown  Screen_Shot_2021-08-29_at_11.16.55_AM.pngScreen_Shot_2021-08-29_at_11.22.03_AM.png
                         Reply  Mon Aug 30 08:04:52 2021, Aidan, DailyProgress, 2um Photodiodes, Monday morning observation of JPL A1 PD cool down  Screen_Shot_2021-08-30_at_8.04.15_AM.png
Message ID: 2639     Entry time: Thu Aug 5 14:29:42 2021     In reply to: 2637     Reply to this: 2650
Author: Stephen 
Type: Laser 
Category: 2um Photodiodes 
Subject: Optical design for 2um PD in new cryo chamber 

Actions on my to-do list, once we are warm and up to air, before Aidan and I are able to run the JPL PD test. This list complements the optical setup tasks and data acquisition setup tasks that are also mentioned by Aidan in this thread.

Update 23 August - using this list to [comment on details] and highlight items which are still outstanding instead of duplicating in a daily progress log entry :)

0. Remove Aaron and Shruti’s Si cantilever clamp, and return to them for safekeeping [I set the clamp aside, need to coordinate return of cantilever to with Aaron!]
1. Solder a connector equivalent to the testpiece (ref. QIL/2465) but with the cryo wire, In-Ag solder on the connector end, and crimped plug pins on the free end. The crimped plug pins interface the with receptacle pins of the existing leads, which are connected to in-vac side of DB9 feedthrough).
 - need to locate some In-Ag solder! [done, and checked connectivity - used standard lead-tin solder following Rana's recommendation (ref. QIL/2418) born out by Koji's testing (ref. QIL/2462 and QIL/2465)]
2. Confirm which PD holder will be most useful for this effort (Koji’s newly machined holders might be useful?) and mount to cold baseplate. If using the tombstone from the IRLabs dewar, which is likely shorter than the beam height, it sounds like we would need to mount it on a pedestal or post. [Aidan used Koji's taller PD mount from the same purchase (ref. QIL/2459). The beam height required no modification, nice!]
3. Mount RTD to PD holder, likely with cryovarnish (unless there is a lucky extra hole for a screw/clamping post). [Radhika and Aidan's G2100807 demo shows the problem with the prior lug, not super stable mounting for the RTD as the same hole is being used to host two screws. Instead, I retrofitted the upper screw from the mount's retaining ring to host one of the trusty spring clamps, see Attachment 1. I checked for clipping or connector interference throughout, and found none.]
4. Mount heater to ___ (TBD, ideally on PD holder but possibly on cold baseplate nearby). [dog-clamped the heater to the baseplate, directly adjacent to the mount - see overview images in Attachment 2 and Attachment 3. We may go through some PID growing pains with this configuration, and we also need to learn whether the 22 W heater power locally applied can overcome the cryocooler's  ~50 W cooling power at our operating temperatures (ref. Radhika's QIL/2585). Might be necessary to intermittently power cycle the cryocooler.]
5. Confirm alignment of shields and PD. [aligned both shields, clamped inner shield, but could reposition if there is an issue.]
6. Verify electrical continuity of PD cable, RTDs, and heater. [note need to add indium and finalize clamping of PD holder, also note routing of pins to be connected to PD connector per Koji's QIL/2605 as described in Attachment 3]
7. Close up (shield lids, chamber lid). [note that in particular, the covering up (with foil sheets) of unused shield apertures is still WIP but wasn't originally mentioned!]
8. (anytime, optional) complete RTD and Heater disconnection junctions with in-vac crimped pins. [done, with mounting isolation achieved by kapton tape as ptfe tubing has not yet arrived. Attachment 4 shows one example, from the inner shield.]
 - crimped plug pins were in 40m lab parts tower, but didn’t see any receptacle pins in the vicinity. [ordered a few hundred new socket pins, I should share some with the 40m parts tower]

Pending Aaron and Shruti’s measurements, it is likely that heater-assisted warmup will occur starting Tuesday, in time for Wednesday/Thursday access. Friday 13th August could be the start of the cool down if everything goes to plan. [Nothing goes well on Friday the 13th. Aaron and Shruti do not need in-vacuum measurements anytime soon. The current plan is for cooldown to begin Tuesday and, if everything goes to plan, we will collect data through the week, then likely swap out the PDs on Monday for another run next week. The next experiment slated for the QIL Vacuum Cryostat is another Si mass radiative cooling run w/ black paint on inner surfaces of inner shield.]
 

Attachment 1: IMG_9685.JPG  1.988 MB  Uploaded Mon Aug 23 22:32:43 2021  | Hide | Hide all
IMG_9685.JPG
Attachment 2: IMG_9687.JPG  3.430 MB  Uploaded Mon Aug 23 23:21:33 2021  | Hide | Hide all
IMG_9687.JPG
Attachment 3: routing_markup_of_IMG_9688.png  11.201 MB  Uploaded Mon Aug 23 23:22:08 2021  | Hide | Hide all
routing_markup_of_IMG_9688.png
Attachment 4: IMG_9682.JPG  2.879 MB  Uploaded Mon Aug 23 23:22:33 2021  | Hide | Hide all
IMG_9682.JPG
ELOG V3.1.3-